OpenStudy (anonymous):

Help with solving this system? :/ A simple walk through would be great -0.2x - 0.3y = 0.2 0.4x - 0.3y = 1.4

OpenStudy (anonymous):

multiply everything by 10

OpenStudy (anonymous):

Why..?

OpenStudy (anonymous):

no decimals

OpenStudy (anonymous):

Oh! Alright let me do that then

OpenStudy (anonymous):

then use the process of elimination to cancel out a variable

OpenStudy (anonymous):

-2x - 3 = 2 && I having trouble with that process. Not understanding it

OpenStudy (anonymous):

This is called a system of equations. You solve it by picking an equation and "isolating" one of the variables. Multiplying everything by 10 first eliminates the decimals and makes it look simpler but is not necessary. but doing so gives: -2x-3y=2 4x-3y=14 solving your first equation for x gets you: x=-1-(3/2)y "(3/2)" is a fraction The next step is to plug what we got for x into the second equation: 4(-1-(3/2)y)-3y=14 Doing this eliminates the x and now you just have one unknown to solve for, y Solving for y gives you: y=-2 plugging that y back into either equation and solving for x gives you x=2

OpenStudy (anonymous):

So is elimination just a bunch of plugging in numbers?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

well somewhat

OpenStudy (anonymous):

Well they accomplish the same thing but to me, if you're learning for the first time, elimination is harder but does the same thing as substitution which is what I just taught you.

OpenStudy (anonymous):

you're trying to cancel one variable out by having the opposite of it

OpenStudy (anonymous):

Alright. Yeah substitution is getting easier for me. But it seems elimination just gets harder and harder..

OpenStudy (anonymous):

In the method of elimination you line the two equations up. In other words, the x term in the first eq. goes above the x term in the 2nd. Same with the y terms and so on. Then you literally subtract the second equation from the first. But, before you do that you need to multiply one of the equations by a number to get either x or y to cancel. In this equation, that wasn't necessary since y is already going to cancel. After either x or y goes away you just solve for whichever one is left. That is pretty much elimination.

OpenStudy (anonymous):

So if the equation was for example (2x + 2y = 2) then you choose which variable you want to cancel?

OpenStudy (anonymous):

well you'd need a 2nd equation for it to be a "system" but yes you sort of choose which one to cancel

OpenStudy (anonymous):

Alright. Starting to make more sense. Thank you (:

OpenStudy (anonymous):

You're welcome (: